# Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions

### v1

@@ 3.4.2 Data preprocessing
Every 16 days, we aggregated all county-level features into a single value to represent agricultural conditions for that period. This method followed the experimental setup from our previous research (Ma et al., 2021b; Wang et al., 2025). We believed that a 16-day interval was sufficient to capture timely changes in weather and environmental conditions.
3.5 Developing the KGML framework
@@ 3.5.1 Pretraining with APSIM field-level dataset
The pretraining process began by using a W2S encoder $f_{W2S}$ to learn the relationship between input weather features $w_{i}$ and predicted soil moisture $\hat{s}_{i}$. This process was formulated as $\hat{s}_{i}=f_{W2S}(w_{i}),\ \text{where}\ w_{i}\in\mathcal{D}_{\text{sim}}$. Then, we used the simulated soil moisture $s_{i}\in\mathcal{D}_{\text{sim}}$ to guide the W2S encoder. The loss function $\mathcal{L}_{\text{SM}}$ was defined as:
$\mathcal{L}_{\text{SM}}=\frac{1}{N}\sum_{i=1}^{N}(s_{i}-\hat{s}_{i})^{2}$ (13)
where $N=\lvert\mathcal{N}\rvert$. Next, we concatenated the predicted soil moisture with weather data and other features to form the features embedding $x_{i}=[w_{i},o_{i},\hat{s}_{i}]$. To predict the final corn yield $\hat{y}_{i}$, we used the attention module $f_{att}$, which took the concatenated features $x_{i}$ as input:
$\hat{y}_{i}=f_{att}(x_{i})$ (14)
The model was trained using the simulated corn yield $y_{i}$ from the dataset $\mathcal{D}_{\text{sim}}$, ensuring that predictions align with PB simulations. The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
@@
Additionally, we incorporated an asymmetric penalty term (Ridnik et al., 2021) controlled by the factor $\lambda>0$, which amplifies the loss when the predicted yield $\hat{y}_{i}$ exceeds the true yield $y_{i}$:
@@
This formulation ensured that the model not only learned accurate yield predictions but also captured the impact of soil moisture variability and drought stress, leading to more reliable and interpretable results.
@@ 3.5.2 Finetuning and testing with GEE county-level dataset
We used the GEE county-level dataset $\mathcal{D}_{\text{train}}$ for finetuning, which contained real-world input features and corresponding USDA NASS yield records. The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were optimized on this dataset. Finally, we evaluated the finetuned model on the testing set $\mathcal{D}_{\text{test}}$, where only the input features were available.
@@ 4.1 Experimental setup
We implemented the deep learning models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011). The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs. For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model. $RMSE$ is a measure of the differences between predicted and actual values in a regression model. It provides an estimate of the standard deviation of the prediction errors. A lower $RMSE$ indicates a better fit of the model to the data. The formula for $RMSE$ is:
@@
$R^{2}$ is a statistical measure that indicates the proportion of the variance in the dependent variable that is predictable from the independent variables. An $R^{2}$ value of 1 indicates perfect prediction, while a value of 0 indicates that the model does not explain any of the variance in the response variable. The formula for $R^{2}$ is:
@@
4.2 Statistical analysis
@@
Before conducting the experiments, we analyzed the statistics of drought, soil moisture, and corn yield. We investigated drought conditions in the United States (Drought.gov, 2024) during August from 2019 to 2023 and present the corresponding maps (Fig. 3). While we focused on August, drought conditions throughout the entire corn growing season from May to October were generally similar. In 2019 (Fig. 3(a)), severe drought was primarily observed in Iowa and Michigan. In 2020 (Fig. 3(b)), severe drought affected Iowa, with North Dakota, South Dakota, Nebraska, and Illinois facing moderate drought. 2021 (Fig. 3(c)) saw worsening conditions, with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought. In 2022 (Fig. 3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought. Finally, in 2023 (Fig. 3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
Figure 3: Drought map of the U.S. in August from (a) 2019 to (e) 2023.
@@
Before studying the impact of soil moisture on corn yield and its prediction, we specified the relationship between soil moisture and drought. Fig. 4 showes the maps of rootzone and surface soil moisture in August from 2019 to 2023. The comparison of the drought maps above with soil moisture data from 2019 to 2023 revealed a strong correlation between drought-affected areas and lower soil moisture levels. We also noticed that rootzone moisture was more abundant than surface moisture and that the two spatial distributions were generally similar.
@@
Figure 4: This figure shows the maps of (1) rootzone and (2) surface soil moisture in August from (a) 2019 to (e) 2023.
@@
Next, we examined the specific impact of drought on corn yield. We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig. 5), where negative values indicate a yield reduction. In 2019 (Fig. 5(a)), drought-affected areas included Iowa and Michigan, however, significant yield reductions were observed only in Michigan. In 2020 (Fig. 5(b)), Iowa faced drought along with a derecho storm (Hosseini et al., 2020), leading to noticeable yield losses. In 2021 (Fig. 5(c)), drought impacted North Dakota, South Dakota, Minnesota, Nebraska, and Iowa, yet significant yield reductions occurred only in North Dakota and South Dakota. In 2022 (Fig. 5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in North Dakota, Kansas, and South Dakota. In 2023 (Fig. 5(e)), although Nebraska and Kansas continued to experience drought, corn yield did not decrease significantly compared to the previous year, likely due to the already reduced yield in 2022. However, Iowa and Missouri experienced some yield losses due to ongoing drought conditions. Thus, we concluded that drought consistently led to a decline in corn yield. We summarized the drought-affected areas that experienced yield reductions in Table 6. In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
@@
4.3 KGML-SM performance comparison with traditional ML models
To validate the superiority of our KGML-SM model, we compared it with four traditional machine learning models: LR, multilayer perceptron (MLP), ridge regression (RR), and random forest (RF). LR (Bishop and Nasrabadi, 2006) is a simple statistical method that models the relationship between a dependent variable and one or more independent variables by fitting a linear equation to the data. MLP (Goodfellow et al., 2016) is a neural network with multiple layers, including an input, hidden, and output layer. It captures non-linear relationships and is widely used for classification and regression. RR (Hoerl and Kennard, 1970) is an extension of linear regression that includes an L2 regularization term to prevent overfitting by penalizing large coefficients. RF (Breiman, 2001) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions to improve accuracy and robustness.
The results showed that our KGML-SM model consistently outperformed the other ML models across all years (Table 7). Based on the results, RF performed the best among traditional ML models, indicating its strong ability to capture complex relationships in the data. RR performed slightly worse than RF, with slightly higher RMSE values, suggesting that regularization helped improve predictions but was not as effective as ensemble learning. LR ranked next, showing higher RMSE values, likely due to its inability to model non-linear relationships effectively. MLP performed the worst, with the highest RMSE values in most years, indicating that it struggled to generalize well, possibly due to overfitting or insufficient training data.
@@ 4.3 KGML-SM performance comparison with traditional ML models
4.4 Ablation study of different compoments in KGML-SM model
To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies. We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025). Next, we incorporated the simulated dataset to evaluate its impact on model performance (Att+sim). We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S). Finally, the impact of the drought-aware loss function was evaluated by comparing the performance of the KGML-SM model with the (Att+sim+W2S) model. We specifically compared the differences among models and their components in Table 8.
@@ 4.4 Ablation study of different compoments in KGML-SM model
Method Attention module Simulated data W2S encoder Drought-aware loss
@@
Through an ablation study of all model components (Table 9), we found that the simulated dataset contributed the most to performance improvement. This indicated that our APSIM field-level dataset effectively captured real-world data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and environmental variables. Additionally, our drought-aware loss function significantly enhanced performance, demonstrating that incorporating prior knowledge of the real-world relationship between drought and corn yield can effectively guide the training and prediction of ML models.
@@ 4.5 Spatial Visualization of model performance
In Fig. 6, we plotted the error maps for the KGML-SM and LR model results from 2019 to 2023. In 2019 (Fig. 6 (a)), the lower part of Michigan shows noticeable overestimation in the attention model, whereas our KGML-SM model performed well. In 2020 (Fig. 6 (b)), although the derecho storm in Iowa also had some impact on our KGML-SM model, we found that the LR model exhibited more severe overestimation. In 2021 (Fig. 6 (c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, is significantly reduced. In 2022 (Fig. 6 (d)), overestimation decreases notably in South Dakota and Kansas, though the effect in Kansas is less pronounced. In 2023 (Fig. 6 (e)), Missouri showed a clear reduction in overestimation, while the effect in Iowa is less noticeable. Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
Figure 6: The error map for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
@@ 4.6 Understanding model bias and prediction-observation alignment
We generated scatter plots for the KGML model and the LR model to analyze their prediction performance (Fig. 7). These plots help visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years. In 2019 (Fig. 7 (a)), the KGML-SM model produced a noticeably narrower distribution, indicating a lower spread in prediction errors. In 2020 (Fig. 7 (b)), The predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the LR model produced more dispersed predictions in high-yield regions. In 2021 (Fig. 7 (c)), the LR model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations. In 2022 (Fig. 7 (d)), the KGML-SM model maintained a narrower and more concentrated prediction distribution, indicating improved robustness. In 2023 (Fig. 7 (e)), the KGML-SM model continued to avoid prediction collapse, demonstrating more stable performance compared to the attention model. This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
Figure 7: The scatter plot for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
@@ 5 Discussion
Specifically, we discussed the role of soil moisture in model prediction from three perspectives:
@@
    How did soil moisture influence model prediction spatially?
@@
    How did soil moisture affect model performance throughout the corn growth season?
@@
    how did soil moisture contribute to model prediction in drought and non-drought regions?
We leveraged the attention mechanism to weight input features, capturing each feature’s contribution to the model, to explore these questions.

### v2

@@ 3.4.2 Data preprocessing
County-level features were aggregated into a single value every 16 days, representing agricultural conditions for that period. This process resulted in a data vector for each feature annually for each county. This aggregation method followed the experimental setup from previous research (Ma et al., 2021b; Wang et al., 2025).
3.5 Developing the KGML-SM framework
@@ 3.5.1 Pretraining with APSIM field-level dataset
The pretraining process began by using a W2S encoder $f_{W2S}$ to learn the relationship between input weather features $\mathbf{w_{i}}$ and predicted soil moisture $\mathbf{\hat{s}_{i}}$. This process was formulated as $\mathbf{\hat{s}_{i}}=f_{W2S}(\mathbf{w_{i}}),\ \text{where}\ \mathbf{w_{i}}\in\mathcal{D}_{\text{field/pretrain}}$. Then, The simulated soil moisture $\mathbf{s_{i}}\in\mathcal{D}_{\text{field/pretrain}}$ is used to guide the W2S encoder. The loss function $\mathcal{L}_{\text{SM}}$ was defined as:
$\mathcal{L}_{\text{SM}}=\frac{1}{N}\sum_{i=1}^{N}(\mathbf{s_{i}}-\mathbf{\hat{s}_{i}})^{2}$ (10)
Next, the predicted soil moisture was concatenated with weather data and other features to form the features embedding $\mathbf{X_{i}}=[\mathbf{w_{i}};\mathbf{o_{i}};\mathbf{\hat{s}_{i}}]$. The final corn yield $\hat{y}_{i}$ was predicted using the attention module $f_{att}$, with the concatenated features $\mathbf{X_{i}}$ as input:
$\hat{y}_{i}=f_{att}(\mathbf{X_{i}})$ (11)
3.5.2 Drought-aware yield prediction loss function
Soil moisture is widely recognized to affect crop yield (Ines et al., 2013; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023). However most prior studies lacked an elegant quantitative approach addressing this issue. This study introduces a loss function adjusting predicted corn yield based on varying soil moisture levels.
The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
@@ 3.5.2 Drought-aware yield prediction loss function
Additionally, there is an asymmetric penalty term (Ridnik et al., 2021) controlled by the factor $\lambda>0$, which amplifies the loss when the predicted yield $\hat{y}_{i}$ exceeds the true yield $y_{i}$:
@@
This whole loss function formulation ensured that the model not only learned accurate yield predictions but also captured the impact of soil moisture variability and drought stress, leading to more reliable and interpretable results.
@@ 3.5.3 Finetuning and testing with GEE county-level dataset
Following the pretraining of the model on the APSIM field-level dataset $\mathcal{D}_{\text{field/pretrain}}$, finetuning using the GEE county-level dataset $\mathcal{D}_{\text{county/finetune}}$ was required to improve the model for the county-level corn yield prediction task. This process involved initially partitioning the data into training, validation, and test sets, followed by predicting corn yield for the target years. Specific details regarding the partition and utilization of these datasets are elaborated upon in the experimental setup (Sec. 4.1). The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were also optimized on the training and validation datasets.
@@ 4.1 Experimental setup
We implemented the DL models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011). The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs. For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model. The formulas for $RMSE$ and $R^{2}$ are:
@@
4.2 Statistical analysis of drought, soil moisture, and corn yield
To study the impact of soil moisture on corn yield in the KGML-SM model and provide interpretability, we first analyzed the statistics of drought, soil moisture, and corn yield. The analysis aims to determine which regions experienced drought and reduced corn yield, and their relationship with soil moisture.
4.2.1 Specifying drought area
@@
We investigated drought conditions in the United States (Drought.gov, 2024) during the corn growth season (June-September) (USDA, 2025) from 2019 to 2023 and present the corresponding maps (Fig. 3). In 2019 (Fig. 3(a)), drought conditions were predominantly observed in Iowa and Michigan. By 2020 (Fig. 3(b)), Iowa experienced severe drought, while moderate conditions affected North Dakota, South Dakota, Nebraska, and Illinois. Conditions deteriorated further in 2021 (Fig. 3(c)), with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought. In 2022 (Fig. 3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought. In 2023 (Fig. 3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
Figure 3: Drought map of the U.S. for (1) June - (4) September from (a) 2019 to (e) 2023.
4.2.2 Soil moisture statistics
@@
Next, we specified the relationship between soil moisture and drought. Fig. 4 shows the maps of average rootzone and surface soil moisture during June-September from 2019 to 2023. The comparison of the drought maps above with soil moisture data from 2019 to 2023 revealed a strong correlation between drought-affected areas and lower soil moisture levels. We also noticed that rootzone moisture was more abundant than surface moisture and that the two spatial distributions were generally similar.
@@
Figure 4: This figure shows the maps of average (1) rootzone and (2) surface soil moisture during June-September from (a) 2019 to (e) 2023.
4.2.3 Drought impact on corn yield
@@
Finally, we examined the specific impact of drought on corn yield. We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig. 5), where negative values indicate a yield reduction. In 2019 (Fig. 5(a)), drought-affected areas included Iowa and Michigan; however, significant yield reductions were observed only in Michigan. In 2020 (Fig. 5(b)), Iowa faced drought along with a derecho storm (Hosseini et al., 2020), leading to noticeable yield losses. In 2021 (Fig. 5(c)), drought impacted North Dakota, South Dakota, Minnesota, Nebraska, and Iowa, yet significant yield reductions occurred only in North Dakota and South Dakota. In 2022 (Fig. 5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in Nebraska, Kansas, and South Dakota. In 2023 (Fig. 5(e)), although Nebraska and Kansas continued to experience drought, corn yield did not decrease significantly compared to the previous year, likely due to the already reduced yield in 2022. However, Iowa and Missouri experienced some yield losses due to ongoing drought conditions. Thus, we concluded that drought consistently led to a decline in corn yield.
@@
4.2.4 Statistical summary of drought-affected states
Based on the above statistical analysis, we summarized the drought-affected areas that experienced yield reductions in Table 7. In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
@@ 4.2.4 Statistical summary of drought-affected states
4.3 Evaluation results
To validate the superiority of our KGML-SM model, we compared it with some commonly used ML models in remote sensing: LR, multilayer perceptron (MLP), ridge regression (RR), and random forest (RF). LR (Bishop and Nasrabadi, 2006) is a simple statistical method that models the relationship between a dependent variable and one or more independent variables by fitting a linear equation to the data. MLP (Goodfellow et al., 2016) is a neural network with multiple layers, including an input, hidden, and output layer. It captures non-linear relationships and is widely used for classification and regression. RR (Hoerl and Kennard, 1970) is an extension of linear regression that includes an L2 regularization term to prevent overfitting by penalizing large coefficients. RF (Breiman, 2001) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions to improve accuracy and robustness.
The results showed that our KGML-SM model consistently outperformed other ML models across all years. As shown in Table 8, RF performed the best among traditional ML models, indicating its strong ability to capture complex relationships in the data (Couronné et al., 2018). RR performed slightly worse than RF, with slightly higher RMSE values, suggesting that regularization helped improve predictions but was not as effective as ensemble learning (Schreiber-Gregory, 2018). LR ranked next, showing higher RMSE values, likely due to its inability to model non-linear relationships effectively (Manual, 2013). MLP performed the worst, with the highest RMSE values in most years, indicating that it struggled to generalize well, possibly due to overfitting or insufficient training data (Caruana et al., 2000).
@@ 4.3 Evaluation results
4.4 Ablation study of different components in KGML-SM model
To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies. We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025). Next, we incorporated the APSIM field-level dataset pretraining to evaluate its impact on model performance (Att+sim). We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S). Finally, the impact of the drought-aware loss function was evaluated by comparing the performance of the KGML-SM model with the (Att+sim+W2S) model. We specifically compared the differences among models and their components in Table 9.
@@ 4.4 Ablation study of different components in KGML-SM model
Method Attention module Field-level data pretraining W2S encoder Drought-aware loss
@@
Through the ablation study of all model components (Table 10), we found that the APSIM field-level dataset pretraining contributed the most to performance improvement. This indicated that our APSIM field-level dataset effectively captured county-level data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and agricultural variables. Additionally, our drought-aware loss function significantly enhanced performance in most years, demonstrating that incorporating prior knowledge of the relationship between drought and corn yield can effectively guide the training and prediction of ML models.
@@ 4.5 Spatial Visualization of model performance
In Fig. 6, we plotted the error maps of the prediction results from 2019 to 2023 for the two best performed models: KGML-SM and RF model. In 2019 (Fig. 6(a)), the lower part of Michigan showed noticeable overestimation in the RF model, whereas our KGML-SM model performed well. In 2020 (Fig. 6(b)), although the derecho storm (Hosseini et al., 2020) in Iowa also had some impact on our KGML-SM model, we found that the RF model exhibited more severe overestimation. In 2021 (Fig. 6(c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, was significantly reduced. In 2022 (Fig. 6(d)), overestimation decreases notably in North Dakota, South Dakota, and Kansas, though the effect in Kansas was less pronounced. In 2023 (Fig. 6(e)), Iowa showed a reduction in overestimation, while the effect in Missouri was less noticeable. Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
Figure 6: The error map for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023, with blue indicating overestimation.
@@ 4.6 Understanding model bias and prediction-observation alignment
To analyze model prediction performance, we generated scatter plots for the KGML-SM model and the RF model (Fig. 7). These plots helped visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years. In 2019 (Fig. 7(a)), the KGML-SM model presented a noticeably narrower distribution, indicating a lower spread in prediction errors. In 2020 (Fig. 7(b)), the predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the RF model produced more dispersed predictions in high-yield regions. In 2021 (Fig. 7(c)), the RF model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations. In 2022 (Fig. 7(d)) and 2023 (Fig. 7(e)), the KGML-SM model maintained a narrower and more concentrated prediction distribution. This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
Figure 7: The scatter plot for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023.
@@ 5 Discussion
In this section, we explore the role of soil moisture in model prediction from four questions:
@@
    (Sec. 5.1) How did soil moisture influence model prediction spatially?
@@
    (Sec. 5.2) How did soil moisture affect model performance throughout the corn growth season?
@@
    (Sec. 5.3) how did soil moisture contribute to model prediction in drought and non-drought regions?
  • 4.
    (Sec. 5.4) How to interpret the observed inaccuracies in corn yield predictions based on soil moisture?